# SAPT: A Shared Attention Framework for Parameter-Efficient Continual Learning of Large Language Models

### v2

@@ Abstract
The continual learning (CL) ability is vital for deploying large language models (LLMs) in the dynamic world. Existing methods devise the learning module to acquire task-specific knowledge with parameter-efficient tuning (PET) block and the selection module to pick out the corresponding one for the testing input, aiming at handling the challenges of catastrophic forgetting and knowledge transfer in CL. However, these methods tend to address only one of the challenges, ignoring the potential of aligning the two modules to effectively address catastrophic forgetting and knowledge transfer simultaneously. To this end, we propose a novel Shared Attention Framework (SAPT), to align the PET learning and selection via the Shared Attentive Learning & Selection module. Extensive Experiments on two CL benchmarks demonstrate the superiority of SAPT. Moreover, SAPT consistently demonstrates its superiority when we scale it to different model sizes (from 770M to 13B), different model architectures (T5 and LLaMA-2) and unseen tasks.
@@
[…]
@@ 1 Introduction
Endowing the continual learning (CL) ability for large language models (LLMs) (Brown et al., 2020; Raffel et al., 2020; Touvron et al., 2023) to learn different tasks sequentially is crucial for their deployment in the real-world, which allows them to dynamically adapt to novel tasks and acquire additional knowledge (Luo et al., 2023; Zhai et al., 2023). However, this scenario presents two significant challenges: (1) Catastrophic Forgetting (CF), referring to the loss of previously acquired knowledge when learning new tasks (McCloskey and Cohen, 1989), and (2) Knowledge Transfer (KT), involving the efficient utilization of knowledge from past tasks to facilitate the learning of new ones (Ke and Liu, 2022).
@@
However, continually updating the SALS leads to the optimal attentive combination only for the newest task, resulting in the forgetting for that of previous ones. Thus, we introduce Attentive Reflection Module (ARM) to help SALS recall what the shared attention operation of samples from previous tasks should be originally performed through generative replay. And the success of ARM offers a new perspective for the utilization of generated pseudo samples instead of just blindly mixing them with samples of new tasks for multi-task training.
@@
  • A novel perspective for the utilization of pseudo samples through generative replay is offered in ARM to assist SALS in navigating inputs to their corresponding PET blocks.
@@ 2.2 Continual Learning for LLMs
are divided into three categories. (1) Rehearsal-based methods introduce the fixed memory to store real samples (Lopez-Paz and Ranzato, 2017; Isele and Cosgun, 2018; Rolnick et al., 2019; de Masson D’Autume et al., 2019) or pseudo-generative examples (Shin et al., 2017; Sun et al., 2019; Su et al., 2019) of previous tasks. (2) Regularization-based methods impose constraints on the loss function to penalize changes regarding the knowledge of previous tasks (Kirkpatrick et al., 2017; Li and Hoiem, 2017; Farajtabar et al., 2020). (3) Parameter isolation methods dynamically expand model capacity or isolate existing model weights to mitigate interference between new and old tasks (Rusu et al., 2016; Fernando et al., 2017).
@@ 4.1 Overview of the Framework
We propose SAPT, a novel framework for the CL of LLMs, offering an effective solution to address the challenges of catastrophic forgetting (CF) and knowledge transfer (KT) simultaneously. The overall architecture of SAPT is illustrated in Figure 2, comprising two key components: (1) Shared Attentive Learning & Selection Module (SALS) and (2) Attentive Replay Module (ARM). In SALS, attentive learning (solid lines) and attentive selection (dashed lines) are aligned through the shared attention operation. Then in ARM, we assist SALS in recalling the exact attentions of inputs from previous tasks with generated pseudo samples.
@@ 4.2 Shared Attentive Learning & Selection Module
In order to transfer the knowledge acquired from previous tasks, when the $t$-th task arrives, parameters of all previous PET blocks $\left\{B_{1},B_{2},\ldots,B_{t-1}\right\}$ and the current one $B_{t}$ are aggregated via weighted combination for the attentive learning of task $\mathcal{T}_{t}$. Specifically, we allocate a key vector $\boldsymbol{k}_{i}$ for each PET block $B_{i}$ ($i\in[1,t])$ and calculate instance-level input-key attentions.111 This process is called shared attention because it will be repeated by the following attentive selection. Such input-key attention ensures the process of attentive learning to be PET-agnostic and compatible with both prompt tuning and LoRA in SAPT.
@@ 5.1.1 Dataset
APnormal-$\uparrow$ F.Ranormal-$\downarrow$ FWTnormal-$\uparrow$ BWTnormal-$\uparrow$ APnormal-$\uparrow$ F.Ranormal-$\downarrow$ FWTnormal-$\uparrow$ BWTnormal-$\uparrow$
@@ 5.1.2 Metrics
Let $a_{i,j}$ be the testing performance (Accuracy for classification task and Rouge-L (Lin, 2004) for others) on the $i$-th task after training on $j$-th task, the metrics for evaluating are:
@@
(2) Forgetting Rate (F.Ra) Chaudhry et al. (2018) measures how much knowledge has been forgotten across the first $\mathcal{T}-1$ tasks, i.e., $F_{\mathcal{T}}=\frac{1}{\mathcal{T}-1}\sum^{\mathcal{T}-1}_{t=1}(a_{t,t}-a_{\mathcal{T},t})$;
@@ 5.3 Implementation Details
SAPT is a model- and PET-agnostic CL method that is compatible with any transformer-based LLM. In our experiments, all methods are performed with instruction tuning Wei et al. (2021); Ouyang et al. (2022) to leverage the task instruction provided in the two benchmarks. To ensure a fair comparison with recent works, we implement SAPT with both prompt tuning and LoRA based on the pre-trained encoder-decoder T5-large model Raffel et al. (2020). We also scale SAPT to the backbone with larger model size (up to 11B and 13B) and the decoder-only LLaMA-2 model Touvron et al. (2023). For the baselines, since they only report the AP metric in their original papers, we carefully re-implement them with their official codes to report metrics of F.Ra, FWT and BWT, providing a thorough insight of how existing methods deal with CF and KT. For more detailed settings, please refer to the Appendix C.
@@
APnormal-$\uparrow$ F.Ranormal-$\downarrow$ FWTnormal-$\uparrow$ BWTnormal-$\uparrow$ APnormal-$\uparrow$ F.Ranormal-$\downarrow$ FWTnormal-$\uparrow$ BWTnormal-$\uparrow$
@@
– DA 44.36 4.16 -2.95 -3.56 67.81 8.24 -8.60 -7.59
Table 2: Results of ablation study on two benchmarks. ARM, Alignment and DA refer to the attentive reflection module, the alignment of the learning and selection in SAPT and shared attentive learning & selection, respectively.
@@ 6.1 Overall Results
Table 1 demonstrates the performance comparison of SAPT and recent PET-based CL baselines on two benchmarks. All results are averaged over the two different orders of each benchmark. Detailed results of each order and each task within a specific order are provided in Appendix D.
@@ 6.3 Ablation Study
After removing the attentive reflection module (“– ARM”, implemented by discarding the $L_{\textrm{KL}}$), the significant decline highlights its crucial role in assisting different input samples to recall the correct shared attention for the corresponding PET blocks they should originally combine. When replacing ARM with naive Replay (“+ Replay”), the decline of F.Ra further verifies our claim that ARM offers a more effective solution to apply pseudo samples.
@@ 6.4 Power of Scale
The results of SAPT and baseline methods on the SuperNI Benchmark based on different sizes of T5 and LLaMA-2 are shown in Figure 4. It can be observed that SAPT is capable of effectively mitigating CF and promoting KT across different model architectures. Moreover, the average performance improves with the enhancement of the model’s basic capabilities (LLaMA-2 > T5). This further demonstrates the generality of our proposed SAPT.
@@
We further select 3 tasks from each one of the above task category to assess the model’s cross-task generalization ability. This is also a crucial dimension for evaluating CL algorithms. Table 3 shows the results. T5-ZS represent the zero-shot approaches for task adaptation, respectively. SAPT yields the best performances, which can be attributed to its superiority in effectively combining acquired knowledge to address novel tasks. This suggests that we should actively promote knowledge transfer between different tasks during the process of CL.
@@ Appendix C Implementation Details
For the generative replay to obtain pseudo samples for our Attentive Replay Module, the prompt length is 300 and is trained for 80 epochs utilizing Adafactor with learning rate of 0.5. And in LoRA, the low-rank $r$ is 8. We train it with AdamW with the learning rate of 0.001 for 5k steps. Batch size is set to 16 for both methods.
Further, we carefully evaluate the official implementations of all baselines, in order to make the comparison as fair as possible. We strictly follow the hyper-parameter settings in their original code, where the prompt size is all set to 10 (except that for LFPT5 of 300) and the LoRA rank is set to 8. If this could not reach the expected performance, we carry out the hyper-parameter search of the learning rate and bachsize for them. Following Sun et al. (2019); Qin and Joty (2022), the volume of replay samples is 0.02 of the original training set for SAPT and all replay baseline methods (Replay and LFPT5). Please refer to Appendix E.2 for deeper analysis for the volume of pseudo samples. All the methods are evaluated for 3 random runs.
@@ Appendix D Fine-grained Results for the Main Experiments
We report the results of each task order on the two benchmark in Table 8 and Table 9. And results of the average performance at each time step is displayed in Figure  7. Overall, the our proposed SAPT demonstrates excellent capabilities in addressing CF and KT.
@@ E.2 Different Volumes and Types of Replayed Samples
In SAPT, the Attentive Reflection Module (ARM) provides a novel perspective for utilizing generated pseudo-data. We conduct additional experiments to analyze the impact of using varying scales of pseudo-data and real data on SAPT and the baseline models Replay and LFPT5. The results are shown in Figure 5. We have the following two observations: (1) Regardless of whether real data or pseudo-data is used, SAPT demonstrates computational efficiency during replay, showing superior performance even with the minimum replay scale [0.02] compared to the maximum replay scale [1] of LFPT5 and Replay. It is worth mentioning that when the replay data volume of Replay is 1, it corresponds to the setting of multi-task learning, which is commonly considered as the upper bound of continual learning. SAPT is able to surpass this upper bound, demonstrating its ability to flexibly handle different inputs, enabling them to be processed by corresponding parameters. (2) For SAPT, there is no significant difference in performance between using real data and pseudo-data. This firstly indicates the reliability of the pseudo-data we generated and the sufficient robustness of our proposed ARM, which can utilize replay data of different qualities to accomplish reflection on shared attention.
@@ 4 T5, LLaMA-2
APnormal-$\uparrow$ F.Ranormal-$\downarrow$ FWTnormal-$\uparrow$ BWTnormal-$\uparrow$ APnormal-$\uparrow$ F.Ranormal-$\downarrow$ FWTnormal-$\uparrow$ BWTnormal-$\uparrow$
@@
APnormal-$\uparrow$ F.Ranormal-$\downarrow$ FWTnormal-$\uparrow$ BWTnormal-$\uparrow$ APnormal-$\uparrow$ F.Ranormal-$\downarrow$ FWTnormal-$\uparrow$ BWTnormal-$\uparrow$
@@
SuperNI task002_quoref_answer_generation Real
@@
[INS] In this task, you’re expected to write answers for questions involving multiple references to the same entity. The answer to the question should be unambiguous and a phrase in the paragraph. Most questions can have only one correct answer. [IN] Passage: In 1899 British President Johnston (now known as John Marshall) is charged with running riots against officers of the Royal Navy. After two days of debriefing, Admiral Sir William Rogers Jr. becomes interested in Robert R̈obJ̈ohnson, another officer on staff, who finds out that Johnson was not present at the rehearsals for the film; however, his father also claims that he has been fired from the navy because Mr Williams had been shot before him. He then went off to see an officer looking after the military. He later admitted himself guilty—for failing which he became involved in the crime? Although no arrest was made during his service they served less than 40 days. He does remain behind but never gets away or die first. He takes control over
task875_emotion_classification Real
@@
[INS] In this task, you are given a sentence containing a particular emotion. You must classify the sentence into one of the six emotions: ’joy’, ’love’, ’anger’, ’fear’, or ’surprise’. [IN] i could clearly feel my adomen muscles contract everytime i cough like some adomen exercise haha and im aching from it now sigh [OUT] sadness
Pesudo
@@
[INS] In this task, you are given a sentence containing monetary hardship. You must classify the sentence into one of the six emotions: ’joy’, ’love’, ’anger’, ’fear’, or’surprise’. [IN] i feel like it is worthwhile to support local artists and so does clay so im amazed by how supportive and passionate artistic writers can be
task1572_samsum_summary Real
@@
[INS] In this task, you are given a conversation, and your task is to generate a summary from the information present in the given conversation. Generate a summary in such a way that the context should be present in the conversation. It should cover the complete context of the conversation. [IN] Mary: Can you pick up the kids?, John: Sure. You’ve already asked me this morning., Mary: Really? I don’t remember. , John: Ok, doesn’t matter. [OUT] John will pick up the kids at Mary’s request.
Pesudo
@@
[INS] In this task, you are given a conversation, and your task is to generate a summary from the information present in the given conversation. Generate a summary in such a way that the context should be present in the conversation. It should cover the complete context of the conversation. [IN] Leda: Ok, so first Homo Faber., Gaby: I’ve already booked the tickets., Ivan: Where exactly?, Leda: Isola San Giorgio., Gaby: They have limited spaces., Ivan: They are always full, but we can ask around it., Gaby: So for us? We won’t make it all day., Marc: OK, what about Giardini or Arsenale? I think we won’t make it to see both. We can ask around., Gaby: I’ll go back!, Ivan: Me too. We could also go different places… There are a lot of satellitanni: What about the transportation? Same ticket
Long Sequence amazon Real
@@
Pesudo
[INS] Given a review text from an e-commerce platform, please analyze the sentiment polarity expressed in the text. Choose one from the following sentiment categories: very negative, negative, neutral, positive and very positive! [IN] Title: Just not as cool as I thought it would be Text: This was just too cool to give up on this movie! Unfortunately, there are way more people with bad taste than Meg McConnell here or even Amy Schumer who make some of her own movies instead of being a good mommy.
imdb Real
@@
Pesudo
[INS] Classify the sentiment polarity expressed in this movie reviews. [IN] Well, I’ve just got back from watching this film for my birthday and it’s been an incredible experience to see how well done that film is! The production values are unbelievable; not only does it look really good, but they have amazing acting: John Wayne makes one of the best actors on his career, and when he plays Jack Ryan (Scott Stierlin) there can be very few performances you would want him to lose weight because he is over 50 pounds. It should all appeal to everybody. There must also explain why such great cinema was important at the time so many years ago. This picture is another highlight which shows how bad James Bond could be like with anything else. We know what we were facing before leaving our country after 24 hours or 48 days…and now we have had some real talent in this film!!
yahoo Real
@@
Pesudo
[INS] I will give you a pair of question and answer, please categorize the topic discussed within. [IN] Question: what is the word "butterfly"? Answer: It means butterfly-like flower that grow in a basket or vase with lilies on it

### v3

@@ Abstract
The continual learning (CL) ability is vital for deploying large language models (LLMs) in the dynamic world. Existing methods devise the learning module to acquire task-specific knowledge with parameter-efficient tuning (PET) block and the selection module to pick out the corresponding one for the testing input, aiming at handling the challenges of catastrophic forgetting and knowledge transfer in CL. However, these methods tend to address only one of the challenges, ignoring the potential of aligning the two modules to effectively address catastrophic forgetting and knowledge transfer simultaneously. To this end, we propose a novel Shared Attention Framework (SAPT), to align the PET learning and selection via the Shared Attentive Learning & Selection module. Extensive experiments on two CL benchmarks demonstrate the superiority of SAPT. Moreover, SAPT consistently demonstrates its superiority when we scale it to different model sizes (from 770M to 13B), different model architectures (T5 and LLaMA-2) and unseen tasks. [^1] Our source code is available at https://github.com/circle-hit/SAPT.
@@
Weixiang Zhao1, Shilong Wang1, Yulin Hu1, Yanyan Zhao1thanks:   Corresponding author, Bing Qin1, Xuanyu Zhang2, Qing Yang2, Dongliang Xu2, Wanxiang Che1 1Harbin Institute of Technology, Harbin, China 2Du Xiaoman (Beijing) Science Technology Co., Ltd. {wxzhao, yyzhao, qinb, car}@ir.hit.edu.cn
@@ 1 Introduction
Endowing the continual learning (CL) ability for large language models (LLMs) (Brown et al., 2020; Raffel et al., 2020; Touvron et al., 2023) to learn different tasks sequentially is crucial for their deployment in the real-world, which allows them to dynamically adapt to novel tasks and acquire additional knowledge (Luo et al., 2023; Zhai et al., 2023; Wu et al., 2024). However, this scenario presents two significant challenges: (1) Catastrophic Forgetting (CF), referring to the loss of previously acquired knowledge when learning new tasks (McCloskey and Cohen, 1989), and (2) Knowledge Transfer (KT), involving the efficient utilization of knowledge from past tasks to facilitate the learning of new ones (Ke and Liu, 2022).
@@
However, continually updating the SALS leads to the optimal attentive combination only for the newest task, resulting in the forgetting for that of previous ones. Thus, we introduce Attentive Reflection Module (ARM) to help SALS recall what the shared attention operation of inputs from previous tasks should be originally performed with pseudo samples. And the success of ARM offers a new perspective for the utilization of generated pseudo samples instead of just blindly mixing them with samples of new tasks for multi-task training.
@@
  • A novel perspective for the utilization of pseudo generated samples is offered in ARM, exhibiting both improved effectiveness and efficiency than naive (generative) replay.
@@ 2.2 Continual Learning for LLMs
are divided into three categories. (1) Rehearsal-based methods introduce the fixed memory to store real samples (Lopez-Paz and Ranzato, 2017; Isele and Cosgun, 2018) or pseudo-generative examples (Shin et al., 2017; Sun et al., 2019) of previous tasks. (2) Regularization-based methods impose constraints on the loss function to penalize changes regarding the knowledge of previous tasks (Kirkpatrick et al., 2017; Li and Hoiem, 2017; Farajtabar et al., 2020; Wu et al., 2022; Chen et al., 2023). (3) Parameter isolation methods dynamically expand model capacity or isolate existing model weights to mitigate interference between new and old tasks (Rusu et al., 2016; Fernando et al., 2017).
@@ 4.1 Overview of the Framework
We propose SAPT, a novel framework for the CL of LLMs, offering an effective solution to address the challenges of catastrophic forgetting (CF) and knowledge transfer (KT) simultaneously. The overall architecture of SAPT is illustrated in Figure 2, comprising two key components: (1) Shared Attentive Learning & Selection Module (SALS) and (2) Attentive Reflection Module (ARM). In SALS, attentive learning (solid lines) and attentive selection (dashed lines) are aligned through the shared attention operation. Then in ARM, we assist SALS in recalling the exact attentions of inputs from previous tasks with generated pseudo samples.
@@ 4.2 Shared Attentive Learning & Selection Module
In order to transfer the knowledge acquired from previous tasks, when the $t$-th task arrives, parameters of all previous PET blocks $\left\{B_{1},B_{2},\ldots,B_{t-1}\right\}$ and the current one $B_{t}$ are aggregated via weighted combination for the attentive learning of task $\mathcal{T}_{t}$. Specifically, we allocate a key vector $\boldsymbol{k}_{i}$ for each PET block $B_{i}$ ($i\in[1,t])$ and calculate instance-level input-key attentions.222 This process is called shared attention because it will be repeated by the following attentive selection. Such input-key attention ensures the process of attentive learning to be PET-agnostic and compatible with both prompt tuning and LoRA in SAPT.
@@ 4.3 Attentive Reflection Module
It is worth to mention that our ARM exhibits both improved effectiveness and efficiency than naive (generative) replay, which is verified by the experimental results in the following Section 6.
@@ 5.1.1 Dataset
[…]
@@ 5.1.2 Metrics
Let $a_{i,j}$ be the testing performance (Accuracy for classification task and Rouge-L (Lin, 2004) for others) on the $j$-th task after training on $i$-th task, the metrics for evaluating are:
@@
(2) Forgetting Rate (F.Ra) Chaudhry et al. (2018) measures how much knowledge has been forgotten across the first $\mathcal{T}-1$ tasks, i.e., $F_{\mathcal{T}}=\frac{1}{\mathcal{T}-1}\sum^{\mathcal{T}-1}_{t=1}(\textrm{max}_{k=i}^{\mathcal{T}-1}a_{k,t}-a_{\mathcal{T},t})$;
@@ 5.3 Implementation Details
SAPT is a model- and PET-agnostic CL method that is compatible with any transformer-based generative LLM. In our experiments, all methods are performed with instruction tuning Wei et al. (2021); Ouyang et al. (2022) to leverage the task instruction provided in the two benchmarks. To ensure a fair comparison with recent works, we implement SAPT with both prompt tuning and LoRA based on the pre-trained encoder-decoder T5-large model Raffel et al. (2020). We also scale SAPT to the backbone with larger model size (up to 11B and 13B) and the decoder-only LLaMA-2 model Touvron et al. (2023). For the baselines, since they only report the AP metric in their original papers, we carefully re-implement them with their official codes to report metrics of F.Ra, FWT and BWT, providing a thorough insight of how existing methods deal with CF and KT. For more detailed settings, please refer to the Appendix C.
@@
[…]
@@
– SA 44.36 4.16 -2.95 -3.56 67.81 8.24 -8.60 -7.59
Table 2: Results of ablation study on two benchmarks. ARM, Alignment and SA refer to the attentive reflection module, the alignment of the learning and selection in SAPT and shared attentive learning & selection, respectively.
@@ 6.1 Overall Results
Table 1 demonstrates the performance comparison of SAPT and recent PET-based continual learning baselines on the SuperNI and Long Sequence benchmarks. All results are averaged over the two different orders of each benchmark. Detailed results of each order and each task within a specific order are provided in Appendix D.
@@ 6.3 Ablation Study
After removing the attentive reflection module (“– ARM”, implemented by discarding the $L_{\textrm{KL}}$), the significant decline highlights its crucial role in assisting different input samples to recall the correct shared attention for the corresponding PET blocks they should originally combine. When replacing ARM with naive Replay (“+ Replay”), the decline of F.Ra further verifies our claim that ARM offers a more effective solution to apply pseudo samples. Please refer to Appendix E.2 for more detailed results and analysis regarding the efficiency of ARM module.
@@ 6.4 Power of Scale
The results of SAPT and baseline methods on the SuperNI Benchmark based on different sizes of T5 and LLaMA-2 are shown in Figure 5. It can be observed that SAPT is capable of effectively mitigating CF and promoting KT across different model architectures. Moreover, the average performance improves with the enhancement of the model’s basic capabilities (LLaMA-2 > T5). This further demonstrates the generality of our proposed SAPT. Please refer to Appendix G for more detailed results.
@@
We further select 3 tasks from each one of the above task category to assess the SAPT’s cross-task generalization ability. This is also a crucial dimension for evaluating CL algorithms. Table 3 shows the results. T5-ZS represent the zero-shot approaches for task adaptation, respectively. SAPT yields the best performances, which can be attributed to its superiority in effectively combining acquired knowledge to address novel tasks. This suggests that we should actively promote knowledge transfer between different tasks during the process of CL.
@@ 8 Limitations
Acknowledgements
We thank the anonymous reviewers for their comments and suggestions. And we thank Libo Qin for the suggestions on the writing of this work.
@@ Appendix C Implementation Details
To obtain pseudo samples for our ARM, the prompt length is 300 and is trained for 80 epochs utilizing Adafactor with learning rate of 0.5. And in LoRA, the low-rank $r$ is 8. We train it with AdamW with the learning rate of 0.001 for 5k steps. Batch size is set to 16 for both methods.
Further, we carefully evaluate the official implementations of all baselines, in order to make the comparison as fair as possible. We strictly follow the hyper-parameter settings in their original code, where the prompt size is all set to 10 (except that for LFPT5 of 300) and the LoRA rank is set to 8. If this could not reach the expected performance, we carry out the hyper-parameter search of the learning rate and batchsize for them. Following Sun et al. (2019); Qin and Joty (2022), the volume of replay samples is 0.02 of the original training set for SAPT and all replay baseline methods (Replay and LFPT5). Please refer to Appendix E.2 for deeper analysis for the volume of pseudo samples. All the methods are evaluated for 3 random runs.
@@ Appendix D Fine-grained Results for the Main Experiments
We report the results of each task order on the two benchmark in Table 8 and Table 9. And results of the average performance at each time step is displayed in Figure  8. Overall, the our proposed SAPT demonstrates excellent capabilities in addressing CF and KT.
@@ E.1 Examples of Pseudo Samples
ARM’s efficiency is demonstrated by its need to generate only the input part of samples, unlike previous generative replay methods (Sun et al., 2019; Qin and Joty, 2022) that required generating complete (input-output) pairs.
@@ E.2 Different Volumes and Types of Replayed Samples
In SAPT, the Attentive Reflection Module (ARM) provides a novel perspective for utilizing generated pseudo-data. We conduct additional experiments to analyze the impact of using varying scales of pseudo-data and real data on SAPT and the baseline models Replay and LFPT5. The results are shown in Figure 6. We have the following two observations that are worth to discuss:
(1) Regardless of whether real data or pseudo-data is used, SAPT demonstrates computational efficiency during replay, showing superior performance even with the minimum replay scale 2% compared to the maximum replay scale 100% of LFPT5 and Replay. It is worth mentioning that when the replay data volume of Replay is 100%, it corresponds to the setting of multi-task learning, which is commonly considered as the upper bound of continual learning. SAPT is able to surpass this upper bound, demonstrating its ability to flexibly handle different inputs, enabling them to be processed by corresponding parameters.
(2) For SAPT, there is no significant difference in performance between using real data and pseudo-data. This firstly indicates the reliability of the pseudo-data we generated and the sufficient robustness of our proposed ARM, which can utilize pseudo data of different qualities to accomplish reflection on shared attention.
@@ Appendix G Scale to LLaMA-2 Model
Selecting O-LoRA as the baseline method for experiments based on LLaMA-2 is because it is the only work among numerous baselines that experimented with LLaMA-2 in the original paper, while other baselines are almost originally implemented with T5 or BERT in their paper. Here we additionally supplement the experimental results of EPI, LFPT5 based on LLaMA2-7B and -13B. Results are shown in Table 10 and Table 11.
@@ 4 T5, LLaMA-2
[…]
@@
[…]
@@
SuperNI Benchmark Long Sequence Benchmark
[…]
Replay 39.48 14.86 0.19 -26.47 71.43 13.64 0.97 -12.73
LFPT5 38.71 16.81 0.32 -15.42 70.31 5.63 0.51 -4.32
EPI - - - - 72.27 5.04 -3.12 -0.50
O-LoRA 24.26 27.56 -6.09 -25.73 54.95 22.36 -8.38 -19.86
SAPT-Prompt 47.39 2.12 0.92 -2.02 77.62 3.29 0.33 -2.98
SAPT-LoRA 56.23 1.07 0.81 -0.65 81.75 2.81 1.09 -2.53
Table 10: The overall results on two continual learning benchmarks with LLaMA-2-7B model. Performance of continual learning (AP), forgetting rate (F.Ra), forward transfer (FWT) and backward transfer (BWT) are reported after training on the last task. All results are averaged over two different orders of each benchmark.
SuperNI Benchmark Long Sequence Benchmark
[…]
Replay 43.99 11.64 0.72 -9.75 76.63 7.92 0.02 -14.86
LFPT5 41.26 14.67 -0.52 -12.31 71.61 6.51 -1.34 -3.78
EPI - - - - 76.66 4.91 -0.09 -1.03
O-LoRA 31.18 24.26 -4.05 -22.64 61.21 19.03 -10.4 -17.54
SAPT-Prompt 52.31 1.57 1.49 -1.41 78.54 3.26 0.14 -2.98
SAPT-LoRA 56.95 1.39 0.81 -0.56 82.32 1.98 0.78 -1.57
Table 11: The overall results on two continual learning benchmarks with LLaMA-2-13B model. Performance of continual learning (AP), forgetting rate (F.Ra), forward transfer (FWT) and backward transfer (BWT) are reported after training on the last task. All results are averaged over two different orders of each benchmark.
@@
Pesudo [INS] In this task, you’re expected to write answers for questions involving multiple references to the same entity. The answer to the question should be unambiguous and a phrase in the paragraph. Most questions can have only one correct answer. [IN] Passage: In 1899 British President Johnston (now known as John Marshall) is charged with running riots against officers of the Royal Navy. After two days of debriefing, Admiral Sir William Rogers Jr. becomes interested in Robert R̈obJ̈ohnson, another officer on staff, who finds out that Johnson was not present at the rehearsals for the film; however, his father also claims that he has been fired from the navy because Mr Williams had been shot before him. He then went off to see an officer looking after the military. He later admitted himself guilty—for failing which he became involved in the crime? Although no arrest was made during his service they served less than 40 days. He does remain behind but never gets away or die first. He takes control over
@@
task875_emotion_classification Real [INS] In this task, you are given a sentence containing a particular emotion. You must classify the sentence into one of the six emotions: ’joy’, ’love’, ’anger’, ’fear’, or ’surprise’. [IN] i could clearly feel my adomen muscles contract everytime i cough like some adomen exercise haha and im aching from it now sigh [OUT] sadness
@@
Pesudo [INS] In this task, you are given a sentence containing monetary hardship. You must classify the sentence into one of the six emotions: ’joy’, ’love’, ’anger’, ’fear’, or’surprise’. [IN] i feel like it is worthwhile to support local artists and so does clay so im amazed by how supportive and passionate artistic writers can be
@@
task1572_samsum_summary Real [INS] In this task, you are given a conversation, and your task is to generate a summary from the information present in the given conversation. Generate a summary in such a way that the context should be present in the conversation. It should cover the complete context of the conversation. [IN] Mary: Can you pick up the kids?, John: Sure. You’ve already asked me this morning., Mary: Really? I don’t remember. , John: Ok, doesn’t matter. [OUT] John will pick up the kids at Mary’s request.
@@
Pesudo [INS] In this task, you are given a conversation, and your task is to generate a summary from the information present in the given conversation. Generate a summary in such a way that the context should be present in the conversation. It should cover the complete context of the conversation. [IN] Leda: Ok, so first Homo Faber., Gaby: I’ve already booked the tickets., Ivan: Where exactly?, Leda: Isola San Giorgio., Gaby: They have limited spaces., Ivan: They are always full, but we can ask around it., Gaby: So for us? We won’t make it all day., Marc: OK, what about Giardini or Arsenale? I think we won’t make it to see both. We can ask around., Gaby: I’ll go back!, Ivan: Me too. We could also go different places… There are a lot of satellitanni: What about the transportation? Same ticket
@@
Pesudo [INS] Given a review text from an e-commerce platform, please analyze the sentiment polarity expressed in the text. Choose one from the following sentiment categories: very negative, negative, neutral, positive and very positive! [IN] Title: Just not as cool as I thought it would be Text: This was just too cool to give up on this movie! Unfortunately, there are way more people with bad taste than Meg McConnell here or even Amy Schumer who make some of her own movies instead of being a good mommy.
@@
Pesudo [INS] Classify the sentiment polarity expressed in this movie reviews. [IN] Well, I’ve just got back from watching this film for my birthday and it’s been an incredible experience to see how well done that film is! The production values are unbelievable; not only does it look really good, but they have amazing acting: John Wayne makes one of the best actors on his career, and when he plays Jack Ryan (Scott Stierlin) there can be very few performances you would want him to lose weight because he is over 50 pounds. It should all appeal to everybody. There must also explain why such great cinema was important at the time so many years ago. This picture is another highlight which shows how bad James Bond could be like with anything else. We know what we were facing before leaving our country after 24 hours or 48 days…and now we have had some real talent in this film!!
@@
Pesudo [INS] I will give you a pair of question and answer, please categorize the topic discussed within. [IN] Question: what is the word "butterfly"? Answer: It means butterfly-like flower that grow in a basket or vase with lilies on it